What is the pressure in a 1.2 L vessel containing 0.32 moles of O2 if the temperature is 23oC (R, the universal gas constant, is 0.082 L atm / mole K.)


a. 0.15 atm

b. 0.50 atm

c. 6.5 atm

d. 8.6 atm


1
Expert's answer
2021-04-02T05:14:46-0400

23oC = 273 + 23 = 296 K

According to the ideal gas law,


"p=\frac{nRT}{V}=\frac{0.32mol\times0.082\frac{L\cdot{atm}}{mol\cdot{K}}\times296K}{1.2L}=6.5atm"


Answer: 6.5 atm (C)